OpenStudy (anonymous):

Need help with Calc :/ s(x)=(3x-9/2x+8)^3 Calculate the derivative of the function.

OpenStudy (mimi_x3):

Use the quotient rule: u = 3x - 9 u' = 3 v = (2x+8)^3 v' = 3(2x+8)^2* 2 = 6(2x+8)^2 Sub it into the quotient rule and solve.

OpenStudy (saifoo.khan):

Mini?

OpenStudy (mimi_x3):

Yes ? Saifi ?

OpenStudy (saifoo.khan):

show me the working please.

OpenStudy (mimi_x3):

Why ? This not your question

OpenStudy (saifoo.khan):

that dosnt matter.

OpenStudy (anonymous):

i have the work

OpenStudy (mimi_x3):

It does matter, I would only show working if the asker ask, NOT you :P

OpenStudy (anonymous):

i got 1134(x-3)^2/(2x+8)^4

OpenStudy (saifoo.khan):

Abnormal people these days

OpenStudy (mimi_x3):

Plagiarism^

OpenStudy (saifoo.khan):

And Lazy people as well.

OpenStudy (mimi_x3):

How am i lazy ?

OpenStudy (saifoo.khan):

coz u arent solving a simple question.. :P

OpenStudy (mimi_x3):

Kendra, i dont think your answer is right.

OpenStudy (mimi_x3):

I don't like giving people the answers these days, I jst show them how to do it hehe If you want the solution, then wait after i finish lunch :P

OpenStudy (saifoo.khan):

lol, you and your lunch.. i know u are having instant noodles like always.. :P i never asked for the answer, i asked u to solve it..

OpenStudy (mimi_x3):

Well, you can solve it yourself.